[Hallc_running] [New Logentry] RC Update

latif at jlab.org latif at jlab.org
Sun Mar 17 17:00:07 EDT 2019

Logentry Text:
- The CSV experiment successfully finished the data taking by 3:30 PM.

- All magnets have been set to zero currents.

- The SHMS has been rotated to 20 deg and HMS to 40 deg.

- The targets are parked now.

- The hall is fully surveyed and in controlled access.

- Detector and beam pipe installation will take place starting from tomorrow.

- Shifts are canceled till Tuesday day shift. Shift takers on Tuesday Swing shift are on standby and should check e-mail prior to coming to the counting-house.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3669226
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190317/627bc73d/attachment.html>

More information about the Hallc_running mailing list